An electrical load operates at 240 V rms The load absorbs an average of 8 Kw at lagging power factor of 0.9 The load impedance is:


 Select one:

 A 5.04 L 45 57 °Ω

 B 3.6 L 60 °Ω

 C 5.76 L 36.87 °Ω

 D 4.32 L 53.13 °Ω

 E 6.48 L 25.84 °Ω


1
Expert's answer
2021-06-04T05:08:02-0400

Average power, P "= 8KW"

But "P=/S/ cos \theta"

Given power factor "cos \theta=0.9"

"=sin \theta=\sqrt{1-cos^2\theta}"

"=sin \theta=\sqrt{1-0.9^2}=\sqrt{0.19}=0.44"

"P=/S/ cos \theta"

"8KW=\/S\/*0.9"

"/S/=\frac{8000}{0.9}=8888.89"

Reactive power, Q "=/S/ sin\theta" VAR

"Q=8888.89(0.44)=3910.72" VAR

"S=P+jQ"

"S=8000+j3910.72"

"S=8+j3.91" KVA

Impedence, Z "\frac{S}{I^2}=4.32 \angle 53.13 °Ω"